OpenStudy (anonymous):

How do you check that T(x,y)=(2x+y,x-y) has additivity or not?

OpenStudy (anonymous):

I know that it equals T(x,y)=T(x)+T(y) but for some reason I'm having a hard time proving it

OpenStudy (zarkon):

you want to try and prove that T((u,v)+(w,z))=T(u,v)+T(w,z)

OpenStudy (zarkon):

T(x,y)=T(x)+T(y) really is nonsense.

OpenStudy (anonymous):

Well then how would I go about doing that? In the book it uses x and y as a reference that's why I got kind of confused

OpenStudy (zarkon):

you want \[T(\alpha+\beta)=T(\alpha)+T(\beta)\] where \[\alpha,\beta \in \mathbb{R}^2\]

OpenStudy (zarkon):

\[T((u,v)+(w,z))=T(u+w,v+z)=(2(u+w)+v+z,u+w-(v+z))\] \[T(u,v)+T(w,z)=(2u+v,u-v)+(2w+z,w-z)=(2u+v+2w+z,u-v+w-z)\] \[=(2(u+w)+v+z,u+w-(v+z))\] notice that they are the same. So additivity holds

OpenStudy (anonymous):

Alright thanks for your help I get it now